1/2

Дмитрий Николаевич Садовников,  Дмитрий Николаевич Садовников,

Из-за острова на стрежень (Стенька Разин)

Sten’ka Razin

tradukita de Tr. Jacob Robbins

1. Из-за острова на стрежень, На простор речной волны, Выплывают расписные, Острогрудые челны.

1. From beyond the wooded island To the river wide and free Proudly sailed the arrow-breasted Ships of Cossack yeomanry.

2. На переднем Стенька Разин, Обнявшись, сидит с княжной, Свадьбу новую справляет, Сам веселый и хмельной.

2. On the first is Stenka Razin With a princess at his side, |: Drunken, holds a marriage revel With his beautiful young bride. :|

3. А она, потупив очи, Ни жива и ни мертва, Молча слушает хмельные Атамановы слова.

3. But behind them rose a whisper, “He has left his sword to woo; |: One short night, and Stenka Razin Has become a woman too!” :|

4. Позади их слышен ропот:

- Нас на бабу променял! Только ночь с ней провозилс Сам наутро бабой стал .

4. Stenka Razin hears the jeering

Of his discontented band,

|: And the lovely Persian princess He has circled with his hand. :|

5. Этот ропот и насмешки Слышит грозный атаман, И могучею рукою Обнял персиянки стан.

5. His black brows have come together As the waves of anger rise, |: And the blood comes rushing swiftly To his piercing, jet-black eyes. :|

6. Брови черные сошлися, Надвигается гроза. Буйной кровью налилис Атамановы глаза.

6. “I will give you all you ask for, Life and heart, and head and hand,” |: Echo rolls the pealing thunder Of his voice across the land. :|

7. “Ничего не пожалею, Буйну голову отдам!” — Раздается голос властный По окрестным берегам.

7. “Volga, Volga, mother Volga, Deep and wide beneath the sun, |: You have never seen a present From the Cossack of the Don. :|

8. ”Волга, Волга, мать родная, Волга, русская река, Не видала ты подарка От донского казака!

8. And that peace might rule as always All my free-born men and brave, |: Volga, Volga, mother Volga, Volga, make this girl a grave.“ :|

…  …

2/2

9. Чтобы не было раздора Между вольными людьми, Волга, Волга, мать родная, На, красавицу возьми!”

9. With a sudden, mighty movement,

Razin lifts the beauty high,

|: And he casts her where the waters

Of the Volga move and sigh. :|

10. Мощным взмахом поднимает Он красавицу княжну И за борт ее бросает В набежавшую волну.

10. Now a silence like the grave sinks

To all those who stand to see,

|: And the battle-hardened Cossacks Sink to weep on bended knee. :|

11. “Что ж вы, братцы, приуныли? Эй, ты, Филька, черт, пляши! Грянем песню удалую На помин ее души!..”

11. “Dance, you fool, and men, make merry!

What has got into your eyes? |: Let us thunder out a chanty Of a place where beauty lies.” :|

12. Из-за острова на стрежень, На простор речной волны, Выплывают расписные Острогрудые челны.

12. From beyond the wooded island

To the river wide and free,

|: Proudly sail the arrow-breasted Ships of Cossack yeomanry. :|

Verkinto de tiu ĉi Rusa poemo estas Дмитрий Николаевич Садовников (Dmitrij Nikolajeviĉ Sadownikow, ∗1847 – †1883).

Arg-478-927 (2010-03-07 18:59:55)

Pri la aŭtoro vidu ankaŭ la informojn sub http: // ru. wikipedia. org , la ruslingva parto de Wikipedia. Informoj pri la kanto, la melodio kaj diversaj versioj de la kanto estas troveblaj en la retejo http: // a-pesni. golosa. info/ popular20/ izzaostrova. htm .

Traduko de la Rusa poemo “Из-за острова на стрежень (Стенька Разин)” de Дмитрий Николаевич Садовников (Dmitrij Nikolajeviĉ Sadownikow, ∗1847 – †1883) en la Anglan de Tr. Jacob Robbins en 1921.

Arg-478-1039 (2010-03-09 17:06:16)

Tiu ĉi angligo troviĝas en http: // ingeb. org/ songs/ stenkara. html .